# Observation of $D^{+}\to\bar{K}_{1}(1270)^{0}\mu^{+}\nu_{\mu}$ and $D^{0}\to K_{1}(1270)^{-}\mu^{+}\nu_{\mu}$

### v1

@@
Observation of $D^{+}\to\bar{K}_{1}(1270)^{0}\mu^{+}\nu_{\mu}$ and $D^{0}\to K_{1}(1270)^{-}\mu^{+}\nu_{\mu}$
@@ Abstract
Experimental studies of the semileptonic decays of charmed mesons are important to deeply understand nonperturbative strong-interaction dynamics in weak decays. The semileptonic $D$ transitions into $\bar{K}_{1}(1270)$ offer an clean window to access the $K_{1}(1270)$ and $K_{1}(1400)$, which are mixtures of the $1^{3}P_{1}$ and $1^{1}P_{1}$ states of $K_{1}$ with mixing angle $\theta_{K_{1}}$. The determination of $\theta_{K_{1}}$ is essential for the theoretical description of $\tau$ Suzuki:1993yc , $B$ Hatanaka:2008xj ; Cheng:2004yj , and $D$ Cheng:2010vk ; Cheng:2003bn decays into axial-vector strange mesons. Throughout this Letter, charged-conjugate modes are implied and $K_{1}$ denotes $K_{1}(1270)$ unless stated otherwise.
@@
Branching fractions (BF) for $D^{+(0)}\to\bar{K}_{1}\ell^{+}\nu_{\ell}(\ell=e$ or $\mu)$ were calculated in the Isgur-Scora-Grinstein-Wise (ISGW) quark model Isgur:1988gb and its update with ISGW2 Scora:1995ty . The BFs for $D^{+}\to\bar{K}_{1}\mu^{+}\nu_{\mu}$ and $D^{0}\to\bar{K}_{1}\mu^{+}\nu_{\mu}$ were predicted to be 0.1% and 0.3%, respectively. However, this model ignores mixing between ${}^{1}\rm P_{1}$ and ${}^{3}\rm P_{1}$ states. Recently, the BFs of $D^{+(0)}\to\bar{K}_{1}\ell^{+}\nu_{\ell}$ were calculated with other theoretical approaches, including three-point QCD sum rules (3PSR) Khosravi:2008jw , the covariant light-front quark model (CLFQM) Cheng:2017pcq , and light-cone QCD sum rules (LCSR) Momeni:2019uag . In general, the predicted BFs range from $10^{-3}$ to $10^{-2}$ Khosravi:2008jw ; Cheng:2017pcq ; Momeni:2019uag ; Bian:2021gwf , and are sensitive to both the amplitude and sign of $\theta_{K_{1}}$. Measurements of the BFs of $D^{+(0)}\to\bar{K}_{1}\mu^{+}\nu_{\mu}$ are important to test these theoretical calculations, to explore the nature of axial-vector strange mesons, and to understand the weak-decay mechanisms of $D$ mesons.
@@
Lepton flavor universality (LFU) is a basic feature of the standard model (SM), in which the couplings between the three families of leptons and the gauge bosons are independent of lepton family. Nevertheless, a few hints of tension between experimental measurements and SM predictions have been reported in some cases, including semileptonic (SL) $B$ decays HFLAV and the Cabibbo-angle anomaly Coutinho:2019aiy ; Crivellin:2020lzu . Intensive tests of LFU in different SL decays of heavy mesons are important to understand these anomalies. Reference Fajfer2015 notes that there may indeed be observable LFU violation effects in the SL decays mediated via $c\to s\ell^{+}\nu_{\ell}$. In the SM, the ratio ${\mathcal{R}}^{\bar{K}_{1}}_{\mu/e}={\mathcal{B}}_{D\to\bar{K}_{1}\mu^{+}\nu_{\mu}}/{\mathcal{B}}_{D\to\bar{K}_{1}e^{+}\nu_{e}}$ is predicted to be $0.95-0.99$ Khosravi:2008jw ; Cheng:2017pcq ; Momeni:2019uag ; Bian:2021gwf . To date, no experimental study of the semimuonic $D^{+(0)}$ decays into axial-vector mesons has been reported, and possible LFU in the SL $D^{+(0)}$ decays into axial-vector mesons has never been tested Ke:2023qzc . Observation of the decays and measurements of the BFs of $D^{+}\to\bar{K}_{1}^{0}\mu^{+}\nu_{\mu}$ and $D^{0}\to K_{1}^{-}\mu^{+}\nu_{\mu}$ offer an important opportunity to test $\mu$-$e$ LFU.
@@
A description of the design and performance of the BESIII detector can be found in Ref. BESIII . The end cap TOF system was upgraded in 2015 using multi-gap resistive plate chamber technology, providing a time resolution of 60 ps BESIII:TOF , which benefits 67% of the data used in this analysis.
Simulated data samples are produced with a geant4-based geant4 Monte Carlo (MC) toolkit including the geometric description of the BESIII detector and the detector response. The simulation includes the beam energy spread and initial state radiation (ISR) in the $e^{+}e^{-}$ annihilations with the generator kkmc kkmc . In the MC simulation, the production of open-charm processes directly produced via $e^{+}e^{-}$ annihilations are modelled with the generator conexc conexc . The ISR production of vector charmonium(-like) states and the continuum processes are incorporated in kkmc kkmc . All particle decays are modelled with evtgen evtgen using BFs either taken from the Particle Data Group pdg2022 , when available, or otherwise estimated with lundcharm lundcharm . Final state radiation from charged final state particles is incorporated using photos photos . The signal MC events of $D^{+(0)}\to K^{-}\pi^{+}\pi^{0(-)}\mu^{+}\nu_{\mu}$ are generated with a special MC generator developed from the amplitude analysis of $D^{+(0)}\to K^{-}\pi^{+}\pi^{0(-)}e^{+}\nu_{e}$ BESIII:k1ff .
@@
For each tag mode, the yield of reconstructed $\bar{D}$ mesons is obtained from an unbinned maximum likelihood fit to its $M_{\rm BC}$ distribution. In the fit, the signal distribution is modeled by the simulated shape convolved with a double-Gaussian resolution function to model the resolution difference between the data and the MC simulation. The combinatorial backgrounds are parameterized by an ARGUS function ARGUS with the endpoint fixed at $E_{\rm beam}=$1.8865 GeV.
@@
Candidates for $D^{+}\to\bar{K}_{1}^{0}\mu^{+}\nu_{\mu}$ or $D^{0}\to K_{1}^{-}\mu^{+}\nu_{\mu}$ are reconstructed from the remaining tracks and showers not used in the $D^{+}$ or $D^{0}$ ST selection. The $\bar{K}_{1}^{0}$ or $K_{1}^{-}$ meson is reconstructed using its dominant decay $\bar{K}_{1}^{0}\to K^{-}\pi^{+}\pi^{0}$ or $K_{1}^{-}\to K^{-}\pi^{+}\pi^{-}$. Candidates for $K^{-}$, $\pi^{\pm}$ and $\pi^{0}$ are selected with the same criteria as those used in the ST selection. To identify muon candidates, we use the information from ${\rm d}E/{\rm d}x$, the time-of-flight system (TOF), and the electromagnetic calorimeter (EMC), and combined confidence levels for the muon, pion, and kaon hypotheses ($CL_{\mu}$, $CL_{\pi}$ and $CL_{K}$) are calculated. Muon candidates are required to satisfy $CL_{\mu}>0.001$, $CL_{\mu}>CL_{K}$ and $CL_{\mu}>CL_{\pi}$. To reduce the background from hadrons, muon candidates are also required to have a deposited energy in the EMC within less than 0.28 GeV. DT candidates must not contain any additional charged tracks, $N^{\rm char}_{\rm extra}=0$; otherwise they are vetoed.
@@
The presence of the undetectable neutrino is inferred via the kinematic quantity $U_{\mathrm{miss}}\equiv E_{\mathrm{miss}}-|\vec{p}_{\mathrm{miss}}|c$, where $E_{\mathrm{miss}}$ and $\vec{p}_{\mathrm{miss}}$ are the missing energy and momentum of the semimuonic candidate, respectively, calculated by $E_{\mathrm{miss}}\equiv E_{\mathrm{beam}}-\Sigma_{j}E_{j}$ and $\vec{p}_{\mathrm{miss}}\equiv\vec{p}_{D}-\Sigma_{j}\vec{p}_{j}$ in the $e^{+}e^{-}$ center-of-mass frame. The index $j$ sums over the $K^{-}$, $\pi^{+}$, $\pi^{0(-)}$ and $\mu^{+}$ of the signal candidate, and $E_{j}$ and $\vec{p}_{j}$ are the energy and momentum of the $j$th particle, respectively. In order to improve the $U_{\mathrm{miss}}$ resolution, the $D$ energy is constrained to the beam energy and $\vec{p}_{D}\equiv-\hat{p}_{\bar{D}}\sqrt{E_{\mathrm{bceam}}^{2}/c^{2}-m_{D}^{2}c^{2}}$, where $\hat{p}_{\bar{D}}$ is the unit vector in the momentum direction of the ST $\bar{D}$, and $m_{D}$ is the nominal $D$ mass pdg2022 .
@@
Figure 1 shows the distributions of $M_{K^{-}\pi^{+}\pi^{0(-)}}$ vs. $U_{\rm miss}$ of the accepted candidates for $D^{+}\to K^{-}\pi^{+}\pi^{0}\mu^{+}\nu_{\mu}$ and $D^{0}\to K^{-}\pi^{+}\pi^{-}\mu^{+}\nu_{\mu}$ in data. Selected $K_{1}$ candidates must have the invariant mass of the $K^{-}\pi^{+}\pi^{0(-)}$ combination within the interval $(1.163,1.343)$ GeV/$c^{2}$, which corresponds to about ±$\pm 1\Gamma$ (decay width) around the $K_{1}$ nominal mass pdg2022 .
@@
Figure 2 shows the distributions of $U_{\rm miss}$ of the accepted candidate events in data. Clear signals around zero are observed. The DT yields are extracted from unbinned extended maximum-likelihood fits to the $U_{\rm miss}$ distributions of data. In the fits, the signal shapes are described by the MC-simulated shape extracted from the signal MC events and convolved with a Gaussian function due to the resolution difference between the data and MC simulation. The combinatorial background shapes are modeled by the MC-simulated shape obtained from the inclusive MC samples. The numbers of signal, the combinatorial background events, and the Gaussian function are free parameters in the fits.
@@
The DT efficiencies, $\varepsilon^{i}_{\rm DT}$, obtained from signal MC samples with the same analysis procedures as for data, are summarized in Table 1 of Ref. supplement . The averaged signal efficiencies in the presence of the ST $\bar{D}$ mesons are $(8.45\pm 0.05)\%$ and $(10.92\pm 0.07)\%$ for $D^{+}\to K^{-}\pi^{+}\pi^{0}\mu^{+}\nu_{\mu}$ and $D^{0}\to K^{-}\pi^{+}\pi^{-}\mu^{+}\nu_{\mu}$, respectively. These efficiencies include correction factors which are discussed later. The final signal efficiency, $\varepsilon_{\rm sig}$, includes the BF of $\pi^{0}\to\gamma\gamma$ pdg2022 . The reliability of the MC simulation was verified by the agreement of the distributions of the momenta and $\cos\theta$ of $K^{-}$, $\pi^{+}$, $\pi^{0(-)}$, and $\mu^{+}$ with the data.
@@
where the first and second uncertainties are statistical and systematic. All related numbers are summarized in Table 1.
@@
The systematic uncertainties in the BF measurement, which are assigned relative to the measured BF, are discussed below. The DT method ensures that most uncertainties arising from the ST selection cancel. The uncertainty associated with the ST yield $N_{\rm ST}^{\rm tot}$, is assigned as 0.3% after varying the signal and background shapes in the $M_{\rm BC}$ fit.
@@
In the studies of $\mu^{+}$ PID efficiencies, the 2D (momentum and $\cos\theta$) PID efficiencies of data and MC simulation of $e^{+}e^{-}\to\gamma\mu^{+}\mu^{-}$ events are re-weighted to match those of $D^{+(0)}\to\bar{K}_{1}\mu^{+}\nu_{\mu}$ decays. The efficiency differences are $-(1.5\pm 0.4)\%$ and $-(2.3\pm 0.4)\%$ for $D^{+}\to\bar{K}_{1}^{0}\mu^{+}\nu_{\mu}$ and $D^{0}\to K_{1}^{-}\mu^{+}\nu_{\mu}$, respectively. The MC efficiencies are then corrected by these differences and used to determine the BFs of the semimuonic decays. After corrections, we assign the uncertainties associated with the $\mu^{+}$ PID to be 0.4% and 0.4%, respectively. The systematic uncertainties due to the combined requirements on $E_{\text{extra~{}}\gamma}^{\rm max},N_{\rm extra}^{\pi^{0}},N^{\rm char}_{\rm extra}$ and, separately, that on $\cos\theta_{\vec{p}_{\mathrm{miss}},\gamma}$ are investigated using the control sample of $D^{+(0)}\to K_{S}^{0}\pi^{0(-)}\mu^{+}\nu_{\mu}$. The corresponding systematic uncertainties are assigned as 3.9% and 4.0% for $D^{+}\to\bar{K}_{1}^{0}\mu^{+}\nu_{\mu}$ respectively; while those for $D^{0}\to K_{1}^{-}\mu^{+}\nu_{\mu}$ are 2.6% and 2.3%, respectively.
@@
To estimate the uncertainties due to the $M_{K^{-}\pi^{+}\pi^{0(-)}}$ requirements, we re-measure the BFs with the nominal $K_{1}$ mass window half width of 0.90 GeV/$c^{2}$ widened to 1.10 GeV/$c^{2}$. The differences between the average BFs and the baseline values are assigned as the systematic uncertainties, which yields 0.6% for both decay modes.
The uncertainties associated with the fits to the $U_{\rm miss}$ distributions are estimated to be $\begin{subarray}{c}+0.5\%\\ -8.8\%\end{subarray}$ ($\begin{subarray}{c}+4.4\%\\ -10.0\%\end{subarray}$) which are dominated by the uncertainty from the background shape. These systematic uncertainties are studied by altering the default MC background shape in three ways. Firstly, alternative MC samples are used to determine the background shape, where the relative fraction of the $q\bar{q}$ background is varied within the uncertainties BESqqbar . Secondly, the BFs of the major $D\bar{D}$ background sources are varied by their uncertainties pdg2022 . Thirdly, the non-$K_{1}^{-}(\bar{K}_{1}^{0})$ sources of $K^{-}\pi^{+}\pi^{-}(K^{-}\pi^{+}\pi^{0})$ ($\begin{subarray}{c}+0.0\%\\ -8.8\%\end{subarray}$ ($\begin{subarray}{c}+0.0\%\\ -9.0\%\end{subarray}$)) are the dominant uncertainties. They are assigned to be the changes of the fitted DT yields after fixing non-resonant components by referring to the non-resonant fraction in $B\to J/\psi\bar{K}\pi\pi$ Belle:2002otd , since the non-resonant fraction in $D^{0(+)}\to\bar{K}\pi e^{+}\nu_{e}$ BESIII:2018jjm ; BESIII:2015hty is close to that in $B^{+}\to J/\psi\bar{K}\pi$ Belle:2010wrf .
@@
To summarize, by analyzing 7.93 $\rm fb^{-1}$ of $e^{+}e^{-}$ collisions data taken at $\sqrt{s}=3.773$ GeV, we report the first observation of $D^{+}\to\bar{K}_{1}^{0}\mu^{+}\nu_{\mu}$ and $D^{0}\to K_{1}^{-}\mu^{+}\nu_{\mu}$ with statistical significances of $12.5\sigma$ and $6.0\sigma$, respectively. Their decay BFs are determined to be ${\mathcal{B}}[D^{+}\to\bar{K}_{1}^{0}\mu^{+}\nu_{\mu}]=(2.36\pm 0.20\begin{subarray}{c}+0.18\\ -0.27\end{subarray}\pm 0.48)\times 10^{-3}$ and ${\mathcal{B}}[D^{0}\to K_{1}^{-}\mu^{+}\nu_{\mu}]=(0.78\pm 0.11\begin{subarray}{c}+0.05\\ -0.09\end{subarray}\pm 0.15)\times 10^{-3}$. Table 3 shows the comparison of the measured BFs with several theoretical calculations. Because the BFs predicted in theory are sensitive to the $K_{1}$ mixing angle, our BFs provides some information on the $K_{1}$ mixing angle.

### v2

# Observation of $D\to\bar{K}_{1}(1270)\mu^{+}\nu_{\mu}$ and test of lepton flavor universality with $D\to\bar{K}_{1}(1270)\ell^{+}\nu_{\ell}$

@@
Observation of $D\to\bar{K}_{1}(1270)\mu^{+}\nu_{\mu}$ and test of lepton flavor universality with $D\to\bar{K}_{1}(1270)\ell^{+}\nu_{\ell}$
@@ Abstract
Experimental studies of the semileptonic (SL) decays of charmed mesons are important to deeply understand nonperturbative strong-interaction dynamics in weak decays. The SL $D$ transitions into $\bar{K}_{1}(1270)$ offer an clean window to access the $K_{1}(1270)$ and $K_{1}(1400)$, which are mixtures of the $1^{3}P_{1}$ and $1^{1}P_{1}$ states of $K_{1}$ with mixing angle $\theta_{K_{1}}$. The determination of $\theta_{K_{1}}$ is essential for the theoretical description of $\tau$ Suzuki:1993yc , $B$ Hatanaka:2008xj ; Cheng:2004yj , and $D$ Cheng:2010vk ; Wang:2020pyy ; Cheng:2003bn decays into axial-vector strange mesons. Throughout this paper, charged-conjugate modes are implied and $K_{1}$ denotes $K_{1}(1270)$ unless stated otherwise.
@@
Branching fractions (BFs) for $D^{+(0)}\to\bar{K}_{1}\ell^{+}\nu_{\ell}(\ell=e$ or $\mu)$ were calculated in the Isgur-Scora-Grinstein-Wise (ISGW) quark model Isgur:1988gb and its update with ISGW2 Scora:1995ty . The BFs for $D^{+}\to\bar{K}_{1}\mu^{+}\nu_{\mu}$ and $D^{0}\to\bar{K}_{1}\mu^{+}\nu_{\mu}$ were predicted to be 0.1% and 0.3%, respectively. However, this model ignores mixing between ${}^{1}\rm P_{1}$ and ${}^{3}\rm P_{1}$ states. Recently, the BFs of $D^{+(0)}\to\bar{K}_{1}\ell^{+}\nu_{\ell}$ were calculated with other theoretical approaches, including three-point quantum chromodynamics (QCD) sum rules (3PSR) Khosravi:2008jw , the covariant light-front quark model (CLFQM) Cheng:2017pcq , and light-cone QCD sum rules (LCSR) Momeni:2019uag . In general, the predicted BFs range from $10^{-3}$ to $10^{-2}$ Khosravi:2008jw ; Cheng:2017pcq ; Momeni:2019uag ; Bian:2021gwf , and are sensitive to both the amplitude and sign of $\theta_{K_{1}}$. Measurements of the BFs of $D^{+(0)}\to\bar{K}_{1}\mu^{+}\nu_{\mu}$ are important to test these theoretical calculations, to explore the nature of axial-vector strange mesons, and to understand the weak-decay mechanisms of $D$ mesons.
@@
Lepton flavor universality (LFU) is a basic feature of the standard model (SM), in which the couplings between the three families of leptons and the gauge bosons are independent of lepton family. Nevertheless, a few hints of tension between experimental measurements and SM predictions have been reported in some cases, including SL $B$ decays HFLAV and the Cabibbo-angle anomaly Coutinho:2019aiy ; Crivellin:2020lzu . Intensive tests of LFU in different SL decays of heavy mesons are important to understand these anomalies. Reference Fajfer2015 notes that there may indeed be observable LFU violation effects in the SL decays mediated via $c\to s\ell^{+}\nu_{\ell}$. In the SM, the ratio ${\mathcal{R}}^{\bar{K}_{1}}_{\mu/e}={\mathcal{B}}_{D\to\bar{K}_{1}\mu^{+}\nu_{\mu}}/{\mathcal{B}}_{D\to\bar{K}_{1}e^{+}\nu_{e}}$ is predicted to be $0.95-0.99$ Khosravi:2008jw ; Cheng:2017pcq ; Momeni:2019uag ; Bian:2021gwf . To date, no experimental study of the semimuonic $D^{+(0)}$ decays into axial-vector mesons has been reported, and LFU in the SL $D^{+(0)}$ decays into axial-vector mesons has never been tested Ke:2023qzc . Observation of the decays and measurements of the BFs of $D^{+}\to\bar{K}_{1}^{0}\mu^{+}\nu_{\mu}$ and $D^{0}\to K_{1}^{-}\mu^{+}\nu_{\mu}$ offer an important opportunity to test $\mu$-$e$ LFU.
@@
A description of the design and performance of the BESIII detector can be found in Ref. BESIII . The end cap TOF system was upgraded in 2015 using multigap resistive plate chamber technology, providing a time resolution of 60 ps BESIII:TOF , which benefits 67% of the data used in this analysis.
Simulated data samples are produced with a geant4-based geant4 Monte Carlo (MC) toolkit including the geometric description of the BESIII detector and the detector response. The simulation includes the beam energy spread and initial state radiation (ISR) in the $e^{+}e^{-}$ annihilations with the generator kkmc kkmc . In the MC simulation, the production of open-charm processes directly produced via $e^{+}e^{-}$ annihilations are modeled with the generator conexc conexc . The ISR production of vector charmonium(like) states and the continuum processes are incorporated in kkmc kkmc . All particle decays are modeled with evtgen evtgen using BFs either taken from the Particle Data Group pdg2022 , when available, or otherwise estimated with lundcharm lundcharm . Final state radiation from charged final state particles is incorporated using photos photos . The signal MC events of $D^{+(0)}\to K^{-}\pi^{+}\pi^{0(-)}\mu^{+}\nu_{\mu}$ are generated with a special MC generator developed from the amplitude analysis of $D^{+(0)}\to K^{-}\pi^{+}\pi^{0(-)}e^{+}\nu_{e}$ BESIII:k1ff .
@@
For each tag mode, the yield of reconstructed $\bar{D}$ mesons is obtained from an unbinned maximum likelihood fit to its $M_{\rm BC}$ distribution. In the fit, the signal distribution is modeled by the simulated shape convolved with a double-Gaussian resolution function to model the resolution difference between the data and the MC simulation. The combinatorial backgrounds are parametrized by an ARGUS function ARGUS with the endpoint fixed at $E_{\rm beam}=$1.8865 GeV.
@@
Candidates for $D^{+}\to\bar{K}_{1}^{0}\mu^{+}\nu_{\mu}$ or $D^{0}\to K_{1}^{-}\mu^{+}\nu_{\mu}$ are reconstructed from the remaining tracks and showers not used in the $D^{-}$ or $\bar{D}^{0}$ ST selection. The $\bar{K}_{1}^{0}$ or $K_{1}^{-}$ meson is reconstructed using its dominant decay $\bar{K}_{1}^{0}\to K^{-}\pi^{+}\pi^{0}$ or $K_{1}^{-}\to K^{-}\pi^{+}\pi^{-}$. Candidates for $K^{-}$, $\pi^{\pm}$ and $\pi^{0}$ are selected with the same criteria as those used in the ST selection. To identify muon candidates, we use the information from ${\rm d}E/{\rm d}x$, the time-of-flight system (TOF), and the electromagnetic calorimeter (EMC), and combined confidence levels for the muon, pion, and kaon hypotheses ($CL_{\mu}$, $CL_{\pi}$ and $CL_{K}$) are calculated. Muon candidates are required to satisfy $CL_{\mu}>0.001$, $CL_{\mu}>CL_{K}$ and $CL_{\mu}>CL_{\pi}$. To reduce the background from hadrons, muon candidates are also required to have a deposited energy in the EMC within less than 0.28 GeV. The DT candidates must not contain any additional charged tracks, $N^{\rm char}_{\rm extra}=0$; otherwise they are vetoed.
@@
The presence of the undetectable neutrino is inferred via the kinematic quantity $U_{\mathrm{miss}}\equiv E_{\mathrm{miss}}-|\vec{p}_{\mathrm{miss}}|c$, where $E_{\mathrm{miss}}$ and $\vec{p}_{\mathrm{miss}}$ are the missing energy and momentum of the semimuonic candidate, respectively, calculated by $E_{\mathrm{miss}}\equiv E_{\mathrm{beam}}-\Sigma_{j}E_{j}$ and $\vec{p}_{\mathrm{miss}}\equiv\vec{p}_{D}-\Sigma_{j}\vec{p}_{j}$ in the $e^{+}e^{-}$ center-of-mass frame. The index $j$ sums over the $K^{-}$, $\pi^{+}$, $\pi^{0(-)}$ and $\mu^{+}$ of the signal candidate, and $E_{j}$ and $\vec{p}_{j}$ are the energy and momentum of the $j$th particle, respectively. In order to improve the $U_{\mathrm{miss}}$ resolution, the $D$ energy is constrained to the beam energy and $\vec{p}_{D}\equiv-\hat{p}_{\bar{D}}\sqrt{E_{\mathrm{beam}}^{2}/c^{2}-m_{D}^{2}c^{2}}$, where $\hat{p}_{\bar{D}}$ is the unit vector in the momentum direction of the ST $\bar{D}$, and $m_{D}$ is the nominal $D$ mass pdg2022 .
@@
Figure 1 shows the distributions of $M_{K^{-}\pi^{+}\pi^{0(-)}}$ vs. $U_{\rm miss}$ of the accepted candidates for $D^{+}\to K^{-}\pi^{+}\pi^{0}\mu^{+}\nu_{\mu}$ and $D^{0}\to K^{-}\pi^{+}\pi^{-}\mu^{+}\nu_{\mu}$ in data. Figure 2 shows the distributions of $M_{K^{-}\pi^{+}\pi^{0(-)}}$ of $D^{+(0)}\to K^{-}\pi^{+}\pi^{0(-)}\mu^{+}\nu_{\mu}$ candidates, in which significant $K_{1}(1270)$ signals can be seen. Selected $K_{1}$ candidates must have $M_{K^{-}\pi^{+}\pi^{0(-)}}$ within the interval $(1.163,1.343)$ GeV/$c^{2}$, which corresponds to about ±$\pm 1\Gamma$ (decay width) around the $K_{1}$ nominal mass pdg2022 .
@@
Fig. 2: The $M_{K^{-}\pi^{+}\pi^{0}}$ (left) and $M_{K^{-}\pi^{+}\pi^{-}}$ (right) distributions of the candidates for $D^{+}\to K^{-}\pi^{+}\pi^{0}\mu^{+}\nu_{\mu}$ and $D^{0}\to K^{-}\pi^{+}\pi^{-}\mu^{+}\nu_{\mu}$. The points with error bars are data, the blue filled histograms are the simulated background, and the red line histograms are the signal MC samples. A requirement of $|U_{\rm miss}|<0.02$ GeV has been applied.
Figure 3 shows the distributions of $U_{\rm miss}$ of the accepted candidate events in data. Clear signals around zero are observed. The DT yields are extracted from unbinned extended maximum-likelihood fits to the $U_{\rm miss}$ distributions of data. In the fits, the signal shapes are described by the MC-simulated shape extracted from the signal MC events and convolved with a Gaussian function due to the resolution difference between the data and MC simulation. The combinatorial background shapes are modeled by the MC-simulated shape obtained from the inclusive MC samples. The signal yield, the combinatorial background yield, and the Gaussian function’s parameters are floated in the fits.
@@
The DT efficiencies, $\varepsilon^{i}_{\rm DT}$, obtained from signal MC samples with the same analysis procedures as for data, are summarized in Table 1 of Ref. supplement . The averaged signal efficiencies in the presence of the ST $\bar{D}$ mesons are $(8.45\pm 0.05)\%$ and $(10.92\pm 0.07)\%$ for $D^{+}\to K^{-}\pi^{+}\pi^{0}\mu^{+}\nu_{\mu}$ and $D^{0}\to K^{-}\pi^{+}\pi^{-}\mu^{+}\nu_{\mu}$, respectively. These efficiencies include correction factors which are discussed later. The final signal efficiency, $\varepsilon_{\rm sig}$, includes the BF of $\pi^{0}\to\gamma\gamma$ pdg2022 . The reliability of the MC simulation has been verified by the agreement of the distributions of the momenta and $\cos\theta$ of $K^{-}$, $\pi^{+}$, $\pi^{0(-)}$, and $\mu^{+}$ with the data.
@@
where the first and second uncertainties are statistical and systematic, respectively. All related numbers are summarized in Table 1.
@@
The systematic uncertainties in the BF measurement, which are assigned relative to the measured BFs, are discussed below. The DT method ensures that most uncertainties arising from the ST selection cancel. The uncertainty associated with the ST yield $N_{\rm ST}^{\rm tot}$, is assigned as 0.3% after varying the signal and background shapes in the $M_{\rm BC}$ fit.
@@
In the studies of $\mu^{+}$ PID efficiencies, the 2D (momentum and $\cos\theta$) PID efficiencies of data and MC simulation of $e^{+}e^{-}\to\gamma\mu^{+}\mu^{-}$ events are reweighted to match those of $D^{+(0)}\to\bar{K}_{1}\mu^{+}\nu_{\mu}$ decays. The efficiency differences are $-(1.5\pm 0.4)\%$ and $-(2.3\pm 0.4)\%$ for $D^{+}\to\bar{K}_{1}^{0}\mu^{+}\nu_{\mu}$ and $D^{0}\to K_{1}^{-}\mu^{+}\nu_{\mu}$, respectively. The MC efficiencies are then corrected by these differences and used to determine the BFs of the semimuonic decays. After corrections, we assign the uncertainties associated with the $\mu^{+}$ PID to be 0.4% and 0.4%, respectively. The systematic uncertainties due to the combined requirements on $E_{\text{extra~{}}\gamma}^{\rm max},N_{\rm extra}^{\pi^{0}},N^{\rm char}_{\rm extra}$ and, separately, that on $\cos\theta_{\vec{p}_{\mathrm{miss}},\gamma}$ are investigated using the control sample of $D^{+(0)}\to K_{S}^{0}\pi^{0(-)}\mu^{+}\nu_{\mu}$. The corresponding systematic uncertainties are assigned as 3.9% and 4.0% for $D^{+}\to\bar{K}_{1}^{0}\mu^{+}\nu_{\mu}$ respectively; while those for $D^{0}\to K_{1}^{-}\mu^{+}\nu_{\mu}$ are 2.6% and 2.3%, respectively.
@@
To estimate the uncertainties due to the $M_{K^{-}\pi^{+}\pi^{0(-)}}$ requirements, we remeasure the BFs with the nominal $K_{1}$ mass window half width of 0.90 GeV/$c^{2}$ widened to 1.10 GeV/$c^{2}$. The differences between the average BFs and the baseline values are assigned as the systematic uncertainties, which yields 0.6% for both decay modes.
The uncertainties associated with the fits to the $U_{\rm miss}$ distributions are estimated to be $\begin{subarray}{c}+0.5\%\\ -8.8\%\end{subarray}$ ($\begin{subarray}{c}+4.4\%\\ -10.0\%\end{subarray}$) which are dominated by the uncertainty from the background shape. These systematic uncertainties are studied by altering the default MC background shape in three ways. First, alternative MC samples are used to determine the background shape, where the relative fraction of the $q\bar{q}$ background is varied within the uncertainties BESqqbar . Second, the BFs of the major $D\bar{D}$ background sources are varied by their uncertainties pdg2022 . Third, the non-$K_{1}^{-}(\bar{K}_{1}^{0})$ sources of $K^{-}\pi^{+}\pi^{-}(K^{-}\pi^{+}\pi^{0})$ ($\begin{subarray}{c}+0.0\%\\ -8.8\%\end{subarray}$ ($\begin{subarray}{c}+0.0\%\\ -9.0\%\end{subarray}$)) are the dominant uncertainties. They are assigned to be the changes of the fitted DT yields after fixing nonresonant components by referring to the nonresonant fraction in $B\to J/\psi\bar{K}\pi\pi$ Belle:2010wrf , since the nonresonant fraction in $D^{0(+)}\to\bar{K}\pi e^{+}\nu_{e}$ BESIII:2018jjm ; BESIII:2015hty is close to that in $B^{+}\to J/\psi\bar{K}\pi$ Belle:2002otd .
@@
To summarize, by analyzing 7.93 $\rm fb^{-1}$ of $e^{+}e^{-}$ collisions data taken at $\sqrt{s}=3.773$ GeV, we report the first observation of $D^{+}\to\bar{K}_{1}^{0}\mu^{+}\nu_{\mu}$ and $D^{0}\to K_{1}^{-}\mu^{+}\nu_{\mu}$ with statistical significances of $12.5\sigma$ and $6.0\sigma$, respectively. Their decay BFs are determined to be ${\mathcal{B}}[D^{+}\to\bar{K}_{1}^{0}\mu^{+}\nu_{\mu}]=(2.36\pm 0.20\begin{subarray}{c}+0.18\\ -0.27\end{subarray}\pm 0.48)\times 10^{-3}$ and ${\mathcal{B}}[D^{0}\to K_{1}^{-}\mu^{+}\nu_{\mu}]=(0.78\pm 0.11\begin{subarray}{c}+0.05\\ -0.09\end{subarray}\pm 0.15)\times 10^{-3}$. Table 3 shows the comparison of the measured BFs with several theoretical calculations. Because the BFs predicted in theory are sensitive to the $K_{1}$ mixing angle, our BFs provides some information on the $K_{1}$ mixing angle. Our BFs disfavor the theoretical calculations from LCSR Momeni:2019uag and CLFQM Bian:2021gwf , which are under the assumption of the mixing angle $\theta_{K_{1}}=-(34\pm 13)^{\circ}$, by more than $5\sigma$.